### v1

# FaaF: Facts as a Function for the evaluation of RAG systems

Abstract
Factual recall from a reference source is crucial for evaluating the performance of Retrieval Augmented Generation (RAG) systems, as it directly probes into the quality of both retrieval and generation. However, it still remains a challenge to perform this evaluation reliably and efficiently. Recent work has focused on fact verification via prompting language model (LM) evaluators, however we demonstrate that these methods are unreliable in the presence of incomplete or inaccurate information. We introduce Facts as a Function (FaaF), a new approach to fact verification that utilizes the function calling abilities of LMs and a framework for RAG factual recall evaluation. FaaF substantially improves the ability of LMs to identify unsupported facts in text with incomplete information whilst improving efficiency and lowering cost by several times, compared to prompt-based approaches.
1 Introduction
The adoption and transformative impact of large language models (LMs) across industries are significantly propelled by their application in knowledge-intensive tasks.
Figure 1: An overview of FaaF, a constructor dynamically creates a function object from a set of fact statements. Function calling allows LMeval to verify all facts within a single call when provided with an input text. FaaF reduces the error rate in identifying unsupported facts by up to 40 percentage points compared to prompting whilst reducing the number of LMeval calls and output tokens by more than 5 times.
Retrieval Augmented Generation (RAG) enables seamlessly the integration of out-of-training knowledge to the LM (Lee et al., 2019). This is achieved by dynamically fetching and incorporating relevant data from a trusted reference source prior LM text generation, resulting in more informed, accurate, and contextually rich outputs. In such applications, factual recall —the extent to which all the information in the reference source that is required to sufficiently answer the question is retrieved and correctly conveyed in the response of the RAG— is a key performance metric because it closely reflects the central purpose of these systems and captures the the performance of retrieval and generation simultaneously.
Although there is a considerable body of work (Chen et al., 2022; Zhang et al., 2023; Gao et al., 2023; Lee et al., 2023; Min et al., 2023; Azaria and Mitchell, 2023; Yuan et al., 2021; Fu et al., 2023; Wang et al., 2023; Kadavath et al., 2022; Manakul et al., 2023) that is focusing at factual precision —the truthfulness of each statement in a generated text— factual recall at least as important for RAG systems, where a response can be factually precise but given the wrong context may be irrelevant to the posed question.
Practical factual recall evaluation in RAG systems is still an open problem because of the following reasons. First, the scenario of automatic verification of truthful, independent statements, given a highly incomplete or inaccurate generated text (simulating recall in a low quality RAG) has not been as explored as precision on LM generated text. Secondly, considering that there can be several facts which need to be verified in a single generated text (E.g. a RAG response to a question), current approaches relying on verification of each fact independently can be prohibitively costly in time and resources. RAG systems include many moving parts (knowledge base, retrieval, prompt formulation, LM) and require substantial tuning (Es et al., 2023) therefore efficiency is a requirement for practical use.
Other recent works (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) attempt to evaluate LMs and RAG systems in terms of factual recall by exact matching of ground truth text in the generated text. While this is a fast and efficient process, it faces serious challenges and limitations in the verification task. As recognised by Cuconasu et al. (2024), exact matching of ground truth text in the generated text is prone to false negatives since the ground truth information might be present in the generated text but phrased differently. Secondly, this approach cannot be applied when the ground truth information longer than a few words since the chances of exact match become too slim.
To address the challenges above, we make the following contributions:
  1. 1.
    We propose an end-to-end factual recall evaluation framework which is tailored to RAG systems. It can be used to (i) create a test dataset and (ii) perform automated factual recall evaluation given a RAG system (or more loosely an LM).
  2. 2.
    We probe into the performance of fact verification formulations in conditions of highly incomplete or inaccurate generated text. To do that, we augment WikiEval [^1] https://huggingface.co/datasets/
    explodinggradients/WikiEval
    (Es et al., 2023) with ground truth facts and human annotation of their validity. WikiEval features question/answer pairs with answers of variable factual quality which enable simulating deficient RAG responses. We find that prompt-based fact verification faces serious challenges in identifying unsupported facts in the presence of inaccurate or incomplete generated text.
  3. 3.
    We introduce Facts as a Function (FaaF), a new fact verification formulation which outperforms fact verification via prompting in all examined conditions and reduces the required number of LM calls and completion tokens by more than 5 times.
  4. 4.
    We open source the factual recall evaluation framework with FaaF as a python package (pip install faaf) and the augmented WikiEval dataset to help the community include factual recall in the RAG optimisation and ultimately build more reliable systems.
2 Related Work
Recently, authors introduced RAGAS (Es et al., 2023), an evaluation framework for RAG systems which is measuring the performance of retrieval and generation without the requirement of ground truth human annotation. Although the authors recognise and are motivated by the need for a practical an efficient RAG evaluation method, RAGAS does not capture the factual aspect of the evaluation, which is a key performance metric and the central focus of RAG’s intended use.
Min et al. (2023) use prompt variations to directly verify individual facts extracted from LM generated biographies. They compare their method with human evaluation and find low error rates when retrieving the ground truth for the evaluated fact.
Zhang et al. (2023) propose self-measuring factuality by the LM via a few-shot prompting method combined by generated facts pertinent to the statement in question. They argue that while leveraging facts from a knowledge base is more dependable, its effectiveness is confined to the scope of the knowledge base and the quality of retrieval. Conversely, self-evaluating with generated facts offers more flexibility but risks introducing inaccuracies.
Li et al. (2023) indicate that LMs have difficulty identifying non-factual information with standard prompting strategies and report improvement using Chain of Thought (CoT).
Azaria and Mitchell (2023) also find that fact verification by prompting insufficient and propose to train a classifier on the hidden layer activations of open source LMs to predict the truthfulness of a generated statements. However, current leading commercial models (GPT-4) are lacking layer activation access and so require alternative methods.
Another approach in the same spirit is to look at the conditional probabilities of each generated token as an indicator of LM confidence and truthfulness of the generated text with the view that low LM confidence is a proxy for incorrect statements (Yuan et al., 2021).
Fu et al. (2023) build on the concept of utilising token probabilities, introducing a self-evaluation framework for LMs. This framework leverages few-shot prompting to evaluate various instructed aspects of LM responses, such as factuallity, fluency, interest, among others.
Manakul et al. (2023) propose SelfCheckGPT which automates the detection of factual errors in LM outputs through statistical analysis of multiple responses to the same prompt, without external knowledge sources. This is again, an expression of the general idea that the probability distribution of the generated response is indicative to the confidence on its truthfulness. So similar to Yuan et al. (2021), SelfCheckGPT makes this assessment post LM-generation by sampling multiple answers on the same prompt thereby removing the requirement of access to the token probabilities or layer weights and making this approach applicable to closed models.
Aly et al. (2021) use a Roberta encoder with a linear layer to learn and predict the fact label given text evidence.
Wang et al. (2023) describe a method where the LM is prompted directly to score an answer’s specific aspect from 0 to 100 or rates it on a 5-star scale, yielding notable results. However, this approach’s effectiveness heavily depends on the prompt’s design.
Zhang et al. (2020) attempt a flexible self-evaluation of generated text using reference answers (BertScore). BertScore calculates a similarity score between tokens in the generated and reference sentences using contextual embeddings. The key benefit being that there is no reliance on exact matching between generated and reference text. Nevertheless, a high semantic score at sentences level does not guarantee factual precision, especially when the information examined is not contextual and explicit (E.g. date).
The work of Zhao et al. (2019) also relies on contextual embeddings but their approach allows for an intentional bias towards precision or recall via reformulating the semantic similarity between generated and reference text as an optimisation problem of finding the minimum effort to transform between the two.
Kadavath et al. (2022) observing that LLMs offer well-calibrated probabilities for multiple-choice questions, effectively transform the validation of model-generated answers into a True/False multiple-choice question. This work offers an important insight in that constraining the the LM response into a binary option, it incites a sufficiently confident response which allows for self-evaluation.
Lastly, recent work on the the factual accuracy of LMs and RAG systems (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) took the approach of using the NaturalQuestions-Open (NQ-open) [^2] https://ai.google.com/research/NaturalQuestions dataset (Kwiatkowski et al., 2019) and calculate accuracy by judging whether any of the ground truth answers (NaturalQuestions annotations) appear in the predicted output via exact matching. NQ-open is a large scale dataset which comprises historical Google search queries and their human-annotated answers sourced from Wikipedia. Even though NQ-open, is valuable for its extensive scope and domain-agnostic nature, fact-verification via exact matches faces serious challenges. As highlighted by Cuconasu et al. (2024), a key issue is accurately determining the truthfulness of answers, especially when dealing with phrases that have the same meaning but (slightly different format or wording) or different date formats. Thus, the necessity for a more advanced analysis of answer variations is recognise and left for future research.
3 Facts as a Function
Facts as a Function (FaaF) introduces a streamlined fact verification method using structured objects for multi-fact assessment. Although in this paper we focus in examining the FaaF approach in the scope of factual recall evaluation in RAG systems, the method itself can equivalently be used for broader fact verification tasks.
Key idea 1: JSON over prompt
We propose that by using the function calling ability of the LM, we enforce a more formal mode of reasoning and communication compared to natural language for a number of reasons. First, by training on code and function definitions the LM learns the independence of the function variables with each other, this results in a clearer individual assessment of each requested variable. In addition, by leveraging function variable metadata and type hints, the LM can be better constrained to the accepted modes of response for each variable. This provides a more consistent guidance compared to instructions given once at the end or the beginning of a prompt. Building on the findings of Kadavath et al. (2022), who established that LMs show well calibrated probabilities when presented with multiple choice questions, we propose that using the function’s type definition of the variables to convey the accepted LM responses is a step further in the same direction, which aids the calibration of the token probabilities. Lastly, due to the strict nature of code syntax compared to natural language, gradients during training are expected to be steeper and thus the LM learns to do a better job at adhering to the expected output and respond with lower stochasticity - which is beneficial in the fact verification use case.
Key idea 2: Generated text as a unit
As discussed above, a function definition can encapsulate a sufficient number of variables to be used by the LM so that they capture all the facts statements which need to be verified. Therefore, we move away from the concept that each fact should be verified individually via a fact-specific prompt and we propose instantiating a function per LM-generated text which needs to be factually assessed. The function can be programatically generated in a manner such that it includes all the individual facts which need to be verified in a given piece of text. In other terms, function calling allows for the fact verification of a long-form text as a unit, with a single LMeval call. This approach results in a reduction of cost and time for fact-verification which is multiple to the number of facts which would otherwise need to be assessed individually, as seen in Chen et al. (2022); Gao et al. (2023); Min et al. (2023); Lee et al. (2023).
Key idea 3: Outsourcing judgement from the LM to the function
Using function definitions to communicate with the LM, enables access to a multitude of tools and further logic processing we can execute on LMeval’s output. This strategy permits us to delegate certain deterministic judgments away from LMeval. We demonstrate this capability by mapping a range of LM responses into a binary format (True / False). In doing so, the calibration of the LM response is enhanced as we provide a more accurate representation of the spectrum of potential outcomes than a simple True/False dichotomy. The underlying intuition is that ultimately, we can afford to ask LMeval simpler and clearer questions which can be answered more reliably and further process the LM output into a final response.
Definition
This work introduces the idea of presenting the facts to the LM as a callable function. Let $S$ be the list of fact statements as strings to verify.
$S=[s_{1},s_{2},\ldots,s_{n}].$
A constructor function $C$ then maps the input list of facts $S$ and control parameters $P$ to an object $O$ with fields $f$ and an object-specific parsing function $G_{O}$.
$C(S,P)\to\left(O\left(f_{1},f_{2},\ldots,f_{n}\right),G_{O}\right).$
Each field from $(f_{1},f_{2},\ldots,f_{n})$ corresponds to a fact statement and is further parameterised by $P$. Control parameters $P$ include the methods and field properties which are injected into resulting object $O$. The object-specific parsing function $G_{O}$ is tasked with parsing the LM response in accordance to the schema and embedded instructions and methods found in $O$. Before passing object $O$ to a language model, we convert it to a JSON representation:
$J(O)\to\text{JSON}_{O}$
Let $M$ be the language model used for fact verification (LMeval). The input of $M$ is a concatenation of $JSON_{O}$, a prompt $q$ which instructs $M$ to utilise $O$ and the input text $x$ which is to be assessed for factual recall with respect to the given facts $S$.
$M(\text{JSON}_{O},q,x)\to o_{x}$
$M$ responds with the output $o_{x}$ which is of string type. Then, the parsing function $G_{O}$ is used to convert the raw response $o_{x}$ to a callable function object $O^{\prime}$ which is an instance with values assigned on its fields.
$G_{O}(o_{x})\to\acute{O}$
Finally, the fact-verification process can be expressed as
$\acute{O}=G_{O}(M(\text{JSON}_{O},q,x)).$
Figure 2: Overview of the factual recall evaluation for RAG. Given a set of ground truth Answers, facts are extracted via LMf. The Hypothesized responses of the RAG (in this instance Ungrounded Answer and Poor Answer) are then tested for recall against the extracted facts.
4 Assessment of fact verification formulations in the RAG setting
Figure 2 outlines the the factual recall evaluation framework which also serves as the experimental setup which we use to compare fact verification formulations with each other. Starting from the ground truth answer containing the desired information to fully answer the posed question, we derive a set of fact statements using a fact-generator LM (LMf). We then use these derived facts to evaluate each of the other answer variants of WikiEval for their factual recall via LMeval. In this manner, each answer is evaluated with respect to the information that is expected from it. It is easy to see how this framework could be applied in the RAG setting where different configurations or model choices impact the quality of the final response to the same input question and knowledge base.
Dataset
In order to probe into performance of automatic fact verification methods we chose to work with the WikiEval dataset (Es et al., 2023) which features three versions of answers, of variable quality, to questions on Wikipedia articles. Specifically, for each question, there is an answer (referred to as ground truth answer in this text for clarity), ungrounded answer and poor answer. All answers have been generated with GPT-3.5-turbo-16k.
The ground truth answer is generated by providing the LM with the correct context from the respective Wikipedia page. The ungrounded answer is generated by not providing any context to the LM. Finally, the poor answer is generated by instructing the LM to give an incomplete answer to the passed question. An example of each answer type can be seen in Figure 2. Generally, ungrounded answer answers feature false, incomplete and redundant information with respect to the ground truth answers. On the other hand, poor answers contain mostly incomplete information compared to ground truth answers i.e. no evidence for support or rejection for ground truth facts.
This dataset enables us to focus on the performance of the fact-verification as a function of the quality and completeness of different answer variants. In that way, we test closely the ability of different fact verification methods to identify unsupported facts when presented with (i) incorrect, (ii) indirectly relevant and (ii) incomplete information with some degree of distinction.
Fact generation
To prepare the WikiEval dataset, we initially generate factual statements that fully capture the information from the ground truth answers, followed by manual annotation of a version of these facts for each answer variant. We use gpt-4-0125 to generate facts from the questions and ground truth answers from the WikiEval dataset (QA pair) using the following prompt:
Convert the given passage into a list of short facts which specifically answer the given question.
Make sure that the facts can be found in the given passage.
The facts should be coherent and succinct sentences with clear and simple syntax.
Do not use pronouns as the subject or object in the syntax of each fact.
The facts should be independent to each other.
Do not create facts from the passage which are not answering the given question.
Add a ”-” before each fact.
Passage: [ground truth answer]
Question: [question]
Fact generation via the prompt above results in a variable number of facts for each ground truth QA pair which depends on the length and information density in the processed ground truth answer. This process yielded 281 individual facts for each type of answer (thus 843 facts in total considering answer, ungrounded answer and poor answer) with an average of 5.6 fact statements generated for every QA pair. The prompt has been designed to ensure that the generated facts are complete sentences, understandable independently of each other or any external references.
Answer Type Human Accuracy
 Answer 100
 Ungrounded Answer 30.24
 Poor Answer 8.53
Table 1: Factual accuracy of the facts derived from the ground truth Answer of WikiEval from human evaluation.
Human fact-verification
We outsource the fact verification of the generated facts against the triplet of answers in WikiEval (answer, ungrounded answer and poor answer) to human evaluators. In this manner we build a ground truth evaluation for each answer type, enabling us to assess the effectiveness of automated fact-verification methods against it. The accuracy from the human fact verification can be seen in Table 1 where the deterioration of the ungrounded answer and poor answer relative to the ground truth answer is evident.
Prompt fact-verification
Following Min et al. (2023), we use a prompt and the respective answer variant as context to verify a single fact at a time with LMeval:
Passage: [answer]
Considering the given passage, the claim [fact] is True or False?
Facts as a function
For each set of fact statements which encapsulate a ground truth answer, we construct a facts-specific object which also contains a parsing method. Since the created object contains all the input facts as fields, we essentially process the set of facts as a unit. A sample of the JSON representation of a function object containing the first fact can be seen bellow:
{’properties’:
    ‘fact_0:{’description’:"Respond
    to the claim in the ’title’ by
    using one of the accepted Enum
    types",
   Ψ’enum’: [’True’, ’False’],
   Ψ’title’: ’It is clear in
    the given passage that Pope
    Benedict XVI became the head
    of the Catholic Church and
    sovereign of the Vatican
    City State on April 19, 2005’
   Ψ’type’: ‘string’},
’required’:
Ψ[‘fact_0’,
 Ψ‘fact_1’
 Ψ …,
 Ψ‘fact_n’]
’title’: ’FactChecker’,
’type’: ‘object’}
Each field includes metadata which can be used to pass instructions, type annotations and the fact statement to be verified itself. In addition to the function object, we pass the following prompt:
Consider the given passage and assign the correct values in the fact checker function.
Passage: [answer]
The answer in the prompt is the input text we want to evaluate against the respective facts which have been previously derived by the ground truth answer in our dataset. After LMeval generates a response, we process it using the tailored parsing method previously created and instantiate the function object that was passed into LMeval, populating it with the generated values in their designated fields.
We test the following configurations:
FaaF(T/F)
A function object with fields which only accept True or False as a response from LMeval.
FaaF(T/F/N)
A function object with fields which only accept True, False or Not clear from the given passage as a response from LMeval. In this scenario, the parser will map Not clear from the given passage to False when processing the LMeval output. This is an example of applying logic post LM generation. The intuition behind this configuration is that the rejection of a claim based on contradicting evidence is conceptually different to the rejection of a claim based on absence of evidence and we help the LMeval’s calibration by providing a clear option for each.
FaaF(T/F)+citation
In this instance we construct a function object with two fields for each input fact. One field for the factual evaluation and one field where we instruct the LMeval to generate an exact excerpt from the input text which directly supports the fact in question (citation field). We place the citation field prior to the factual evaluation field so that LMeval is made to first try and find a supporting citation from the input text before verifying the fact that is being assessed. Similarly, the intuition here is that by asking LMeval to search and retrieve a specific citation from the input text which supports a specific fact, it will result in a better calibrated verification of the respective fact statement.
FaaF(T/F/N)+citation
In this configuration we combine the two approaches outlined above to explore their combined effect. In detail, we construct the function object to include citation fields and factual evaluation fields for each fact and we define True, False or Not clear from the given passage as accepted responses.
Language models (LMeval)
We use the latest commercially available model which support function calling gpt-4-0125. A comprehensive assessment of open source models which support function calling is left for future study as there are none available at this time but we will extend this study to more models as they become available. Also, the focus of this work is on the different formulation of facts during verification and not the assessment of specific LMs.
Metrics
Error rate
We use Error Rate (ER) between the human fact-verification and the fact verification formulation as the main indicator of their reliability.
F1micro
We also use F1micro score (F1m) as defined in Min et al. (2023) to measure the successful identification of unsupported facts and probe further into the individual fact verification. It should be noted that F1 scores explicitly depend on the class ratio (T/F) via precision and recall. For that reason F1m scores should be compared across fact verification approaches in the same answer category (where the T/F ratio is preserved) in Table 2, and not across answer categories.
Facts Formulation (ground truth) Answer Ungrounded Answer Poor Answer
N/A ER F1m N/A ER F1m N/A ER F1m
Prompt(T/F) 0/281 2.1 $*{}^{*}$ 4/281 26.7 77.6 3/281 50.8 61.8
FaaF(T/F) 0/281 1.7 $*{}^{*}$ 0/281 24.9 79.6 0/281 9.9 94.2
FaaF(T/F/N) 0/281 1.4 $*{}^{*}$ 0/281 18.1 85.7 0/281 7.4 95.7
FaaF(T/F)+citation 0/281 0.7 $*{}^{*}$ 0/281 22 82.5 0/281 7.8 95.6
FaaF(T/F/N)+citation 0/281 0.3 $*{}^{*}$ 0/281 23.1 81.7 0/281 7.4 95.8
Table 2: Results on Error Rate (ER) along with F1micro estimated by each fact formulation. F1micro measures accuracy and recall in predicting the false facts. $*{}^{*}$1micro is 0 in the Ground Truth Answer cases because there are no False facts in the ground truth evaluation in this category. T/F indicates True/False and T/F/N are True/False/Not clear from the given passage respectively. N/A indicates the Not answered i.e. the number of facts that the LM did not respond on their validity.
5 Results
Table 2 presents the non-response rate (N/A), error rate and F1m from the fact verification formulations across the answer categories which are examined.
Prompting for verification is not reliable in cases of incorrect or incomplete information. Although in the case of ground truth answers Prompt(T/F) shows 2.1%ER (in line with what is reported in Min et al. (2023) when retrieval is enabled), we see a sharp rise when we try to verify facts in text from the ungrounded answers (26.7%) and a further deterioration with poor answers (50.8%). Such performance indicates that prompting is not suitable for fact verification in text with unknown information quality and completeness.
FaaF outperforms zero shot prompt in all cases. Formulations which leverage function calling demonstrate a notable improvement in ER and F1micro over all cases (answer type, FaaF configuration). The improvement is most pronounced in the poor answer category where ER is reduced from 50.8% (prompting) to less than 10% with FaaF methods denoting a paradigm shift in LMeval’s ability to identify missing information in the input text. The milder improvement seen in the ungrounded answer category suggests that elucidating the fact verification in the presence of incorrect and indirectly relevant information is more challenging for LMeval. Lastly, in the case of prompting, LMeval failed to respond on 4/281 of the examined facts for the ungrounded answer and 3/281 for the poor answer. Conversely, LMeval returned a valid response for all facts, in all FaaF configurations and answer types.
LMs tend to overestimate fact truthfulness overall. The human evaluation of factual accuracy in ground truth answers is 100% i.e. every fact is True (Table 1). This coincides with the lowest ER scores in Table 2, irrespective of the fact verification approach.
Rejecting facts is more difficult in the presence of tangentially relevant information when using FaaF. It is evident from Table 2 that FaaF approaches struggle more with ungrounded answers than with poor answers. It should be noted that this observation is reversed when using prompt(T/F), where the ER is higher in poor answers, due to the significant overestimation of truthfulness this approach shows in the absence of supporting information. The key difference between the two answer types is that ungrounded answers typically contain incorrect or tangentially relevant information to the facts that are examined whilst poor answers usually lack information altogether. This is because ungrounded answers are generated by an LM without any context thus, depending on the popularity and prevalence of the theme of the question, a degree of factual hallucination can be expected.
Consider the following example from WikiEval:
Question:
What factors contributed to the Sri Lankan economic crisis?
Ground truth answer:
The Sri Lankan economic crisis was caused by multiple compounding factors. These include tax cuts and money creation by the government, which led to budget deficits and a decline in government revenue. The government’s decision to shift to organic or biological farming also played a role. Additionally, the crisis was exacerbated by the 2019 Sri Lanka Easter bombings and the impact of the COVID-19 pandemic. The country’s high external debt, with a significant increase in foreign debt in recent years, also contributed to the crisis. The combination of these factors resulted in unprecedented levels of inflation, near-depletion of foreign exchange reserves, shortages of medical supplies, and an increase in prices of basic commodities.
Ungrounded answer:
There were several factors that contributed to the Sri Lankan economic crisis:
  1. 1.
    High levels of government debt: Sri Lanka had accumulated a significant amount of government debt, both domestic and foreign. This debt burden became unsustainable, leading to a crisis in public finances.
  2. 2.
    Fiscal mismanagement: The government had been running large budget deficits for several years, spending more than it was earning. This led to a widening fiscal deficit and increased borrowing to finance government expenditure.
  3. 3.
    Political instability: Sri Lanka experienced political instability during this period, with frequent changes in government and policy uncertainty. This created an unfavourable business environment and deterred foreign investment.
  4. 4.
    Declining exports: Sri Lanka’s export sector, particularly garments and textiles, faced challenges due to global economic slowdowns and increased competition from other countries. This led to a decline in export earnings, which further strained the country’s external finances.
  5. 5.
    High inflation: Sri Lanka experienced high inflation during this period, which eroded the purchasing power of the population and increased the cost of living. This put additional pressure on the economy and contributed to the crisis.
  6. 6.
    External shocks: Sri Lanka was also affected by external shocks, such as rising oil prices and global financial market volatility. These factors further exacerbated the economic challenges faced by the country.
  7. 7.
    Structural issues: Sri Lanka had several structural issues that hindered its economic growth, such as a lack of diversification in the economy, inadequate infrastructure, and low productivity levels. These issues made it difficult for the country to sustain economic growth and overcome the crisis.
Poor answer:
There were several factors that contributed to the Sri Lankan economic crisis.Sri Lanka is known for its beautiful beaches and rich cultural heritage.
Extracted facts(using question and ground truth Answer):
  • Tax cuts and money creation by the government contributed to the Sri Lankan economic crisis.
  • Budget deficits and a decline in government revenue were factors in the Sri Lankan economic crisis.
  • The governments decision to shift to organic or biological farming played a role in the crisis.
  • The 2019 Sri Lanka Easter bombings exacerbated the economic crisis.
  • The impact of the COVID 19 pandemic contributed to the Sri Lankan economic crisis.
  • High external debt, with a significant increase in foreign debt in recent years, also contributed to the crisis.
Although the ungrounded answer is quite verbose and has several mentions and indirect references of the extracted facts, it fails to capture with clarity the information from the ground truth answer which would allow for their confident verification. Meanwhile, the poor answer caries no useful information in this instance.
In this scenario LMeval has a higher risk of a misjudgement in the ungrounded answer than the poor answer. This seems coherent intuitively since rejecting a claim in the presence of relevant information is a more demanding and complex task which requires deeper interpretation of the language than when there is no relevant information.
Figure 3: LMeval call count for a full evaluation of WikiEval. FaaF formulations result in more than five times less LM calls.
Facts formulation Prompt Tokens Completion Tokens Total Tokens
Prompt(T/F) 146,276 26,545 172,821
FaaF(T/F) 69,797 5,658 75,455
FaaF(T/F/N) 77,384 7,113 84,497
FaaF(T/F)+citation 122,600 20,596 143,196
FaaF(T/F/N)+citation 130,187 21,889 152,076
Table 3: LMeval token count for the tested fact formulations, for the factual recall evaluation of WikiEval dataset.
Providing a “not clear” option helps. We observe a reduction of the error rate and corresponding increase in F1m when we add the option for LMeval to respond with Not clear from the given passage which is mapped to False as a post-generation step in the parsing function. The helpful mechanism in this instance is that we provide a needed third option to LMeval when the token probability distribution between True / False is not clearly indicating one over the other. Rejecting a statement due to conflicting evidence and due to lack of evidence are both as valid reason as they are distinct to each other. Using False to capture both rejection scenarios has proved to lead to more false positives than providing LMeval with the option to distinguish between them.
Citations help only in the presence of correct and clear information. The positive impact of adding citations is most evident in the ground truth answer category where the provided text always contain the required evidence to support the facts. In this instance, asking for text evidence from LMeval results in avoiding some false negative verifications. The beneficial mechanism is associated with inserting a preliminary step to the fact verification process - of the explicit use of evidence from the input text. This aligns with the findings reported in Wei et al. (2023) regarding the chain-of-thought method.
Interestingly, for the other answer categories, the citation benefit becomes less clear and even reversed. ER is relatively stable in poor answers and it is seen to increase in the case of ungrounded answer. In detail, we notice two conflicting trends: firstly, citations can prevent false positives by highlighting the absence of supporting text for a given fact statement. Conversely, in a few instances, the lack of explicit citations leads to false negatives, as LMeval fails to recognize a fact as true, even though a holistic human interpretation of the answer would deem it accurate.
LMeval usage efficiency
Beyond the improvements in ER and F1m, another advantage in encapsulating multiple facts into a function object is the that of efficiency. Figure 3 shows reduction in the number of calls to LMeval required for the factual evaluation of WikiEval by more than five times. The number of calls to LMeval required for FaaF corresponds to the number of texts to be evaluated, whereas, in prompt-based fact verification, it matches the number of fact statements.
Table 3 indicates that FaaF(T/F) uses half the prompt tokens and less than one-fifth of the completion tokens required by prompt-based fact verification for a complete evaluation on WikiEval(all answer types). Particularly in the setting of RAG systems optimisation, the efficiency and scalability of an evaluation framework is of high importance since it is proportionally related to the number of RAG configurations which can be tested.
Including citations and the option Not clear from the given passage as acceptable responses in FaaF affects the token count, significantly increasing it. However, it remains below the token requirements of prompt-based fact verification. In essence, FaaF makes possible fact verification with citations at a lower token cost than prompt-based verification without citations.
6 Conclusions & future work
We show that prompt-based fact verification is prone to overestimating the truthfulness of fact statements in texts with inaccurate and/or missing information. In particular the error rate of prompting can reach up to 50% when the text under review is significantly lacking in information. In such challenging situations, presenting the facts as a function to the LM significantly enhances both the error rate and the F1m measure of the verification process. Using FaaF, we observe that texts with tangentially relevant and inaccurate information’s are more likely to cause false positives than texts with missing or incomplete information. By testing various configurations with FaaF, we find that by including a “not clear” option to the True/False dichotomy helps overall. The impact of asking for citations before fact verification is sensitive to the quality and coverage of information in the examined text. While beneficial when fact-supportive information is present or clearly absent, this approach can lead to false negatives if the text can support the fact as a whole but lacks direct, extractable citations.
Finally we report significant cost and time efficiency improvements between prompting and function based fact representations. Generally, using FaaF leads to a reduction in both the number of calls to the LM and the number of tokens needed for fact verification by a multiple-factor.
Limitations
While the advantages of using function calls for fact verification are significant, there is certain variance expected in the results due to the stochastic nature of LM generation. Further extensive testing is necessary to solidify the results presented, especially probing into the influence of citations in the performance of verification using function objects. Although, the WikiEval dataset has been instrumental in providing a convenient way to compare fact verification performance across various text qualities, it is relatively small, comprising only 50 question/answer pairs.
Additionally, the results shown with FaaF are sensitive to the instructions passed in the function object’s metadata, which highlights the need for additional research into different FaaF configurations. Other open questions include the maximum number of fact statements and the maximum permissible length for a fact that can be incorporated into a function object and whether token count is the sole limitation or if there are performance implications as well.
Acknowledgements
This research was supported by IMMO Capital, London UK.